### v1

@@ Abstract
This paper examines the identification and estimation of treatment effects in staggered adoption designs – a common extension of the canonical Difference-in-Differences (DiD) model to multiple groups and time-periods – in the presence of (time varying) misclassification of the treatment status as well as of anticipation. We demonstrate that standard estimators are biased with respect to commonly used causal parameters of interest under such forms of misspecification. To address this issue, we provide modified estimators that recover the Average Treatment Effect of observed and true switching units, respectively. Additionally, we suggest a testing procedure aimed at detecting the timing and extent of misclassification and anticipation effects. We illustrate the proposed methods with an application to the effects of an anti-cheating policy on school mean test scores in high stakes national exams in Indonesia.
@@
Keywords: Measurement Error, Causal Inference, Moment Equality Tests, Misspecification Tests
@@ 1 Introduction
The Difference-in-Differences (DiD) methodology is arguably one of the most popular research designs in applied economics, offering a simple and intuitive framework to estimate causal effects in observational settings. It serves as a preferred tool for policy evaluations, labor market studies, and other areas where randomized controlled trials are impractical. Staggered adoption designs naturally extend the canonical two-by-two DiD framework to settings where multiple groups adopt treatments at different times. This structure enables researchers to exploit variation in treatment timing and group-level adoption, aggregate treatment effects across groups, and accommodate heterogeneity in treatment effects. [^1] To illustrate the prevalence of staggered adoption designs, Baker et al. (2022) report that more than 54% of all DiD studies published in the leading Finance and Accounting journals between 2000 and 2019 (744 studies in total) were based on staggered adoption designs.
@@
Misclassification of the treatment status arises when the recorded treatment status does not coincide with the actual treatment status. This can occur due to data limitations, such as when the exact timing of treatment is not observed (e.g., Bindler and Hjalmarsson, 2018), or due to reporting and classification errors when treatment status can only be inferred indirectly (see Goldschmidt and Schmieder, 2017). Anticipation effects, on the other hand, occur when individuals or units are either exposed to treatment prior to the recorded treatment date or adjust their behavior in anticipation of future treatment, thereby contaminating pre-treatment periods (see, e.g., Barrios, 2022; Bismarck-Osten et al., 2022; Berkhout et al., 2024). Both issues are particularly problematic in staggered adoption designs, where treatment assignment evolves over time and across groups, amplifying the risk of “forbidden comparisons” and leading to invalid inference, even when estimands appropriate for staggered adoption settings are targeted.
@@
As a second contribution, we define alternative estimators that allow to recover the aforementioned causal parameters in staggered adoption designs. In particular, we introduce a bias-corrected version of the staggered adoption estimator proposed by De Chaisemartin and D’Haultefoeuille (2020), which recovers the ATE of observed switching-to-treatment units. Based on knowledge of the timing and extent of misspecification, the estimator adjusts the comparisons conducted by the original estimator of De Chaisemartin and D’Haultefoeuille (2020) by using groups unaffected by misspecification. In contrast, for the ATE of actual switching-to-treatment cells, additional steps are required. Since the actual timing of treatment adoption is unobserved at the unit level, we exploit a homogeneity assumption on the misclassification (anticipation) probability, which restricts heterogeneity across units but not over time. This allows us to consistently estimate the fraction of actual switching-to-treatment units by extending the Parallel Trends (PT) assumption to the (potential) outcomes under treatment. In a subsequent step, we construct an estimator for the ATE of actual switching-to-treatment cells and derive its asymptotic properties.
As a third contribution, we propose a two-step testing procedure to detect the presence, extent, and timing of misspecification. These tests provide guidance on whether and when misclassification or anticipation occurs, which is required by the estimators introduced in the previous section. In particular, the corrections proposed in this paper rely on common PT assumptions. However, distinguishing between bias arising from misspecification and bias due to PT violations in pre-treatment periods is inherently challenging. To address this, we propose a two-step moment equality testing framework that combines a pre-trend test with a test for misspecification. The first step aims to maximize confidence in the PT assumption by examining pre-treatment evidence for its validity, comparing not-yet-switched comparison groups with never treated groups. Conditional on passing this stage, the second step tests for misclassification or anticipation by comparing observed switcher groups to not-yet-switched groups using lagged periods. We implement these tests as moment equality tests and derive their asymptotic properties.
@@
To illustrate the finite sample performance of both the estimators and the tests, we carry out a small-scale Monte Carlo simulation. The findings suggest a good performance even in cases of misspecification because of misclassification or anticipation. We then apply these methods to assess the effects and the presence of anticipation in a study by Berkhout et al. (2024) on the effects of a new computer based testing system on school mean test scores in Indonesia between 2015 to 2019. The testing system, which was voluntary and intended to curb cheating behavior, was reported to have led schools to anticipate measures that limit cheating. Our estimates for the ATE of observed switching-to-treatment units confirm this and align with Berkhout et al. (2024), while the estimates for the ATE of actual switching-to-treatment units are slightly larger in magnitude. Moreover, we estimate the fraction of anticipating schools to be around 15% in 2018, and up to 30% in 2019. Finally, the misspecification tests provide evidence for anticipation effects, though they also suggest some violation of parallel pre-trends in periods just prior to that anticipation for at least one cohort.
@@
By providing a framework for inference in staggered adoption designs under misspecification, this paper contributes to the growing literature on the methodological underpinnings of DiD. In particular, the paper builds on three strands of the econometric literature. First, it contributes to the growing body of research on DiD methods tailored to staggered adoption designs, including recent advancements that address the challenges of heterogeneous treatment effects and dynamic treatment timing (e.g., De Chaisemartin and D’Haultefoeuille, 2020; Callaway and Sant’Anna, 2021; Borusyak et al., 2024). Second, the paper intersects with a broader literature on the misclassification of treatment status, which is well known to compromise the validity of inference procedures (e.g., Lewbel, 2007b; Ura, 2018; Denteh and Kedagni, 2022). Unlike other approaches in the DiD literature that rely on proxy variables (e.g., Botosaru and Gutierrez, 2018) or exclusion restrictions (Negi and Negi, 2025), our framework does not require such external information. Third, the paper relates to recent contributions on pre-testing of Parallel Trends and its implications for causal inference depending on the test outcome (e.g., Rambachan and Roth, 2023; De Chaisemartin and D’Haultefoeuille, 2024).
@@ 2 Setup
[…]
@@
Similar to De Chaisemartin and D’Haultefoeuille (2020), we replace the independent groups assumption with mean independence, since the presence of stable groups (as defined below) inherently violates independence across groups (specifically, observed treatment status cannot be independent across groups in this case). Moreover, our sharp design assumption (Assumption 2) is stated with respect to the observed treatment status $D_{i,g,t}$, thus explicitly accommodating scenarios in which researchers mistakenly assume a sharp design based solely on observed data. Importantly, we allow for violations of the sharp design at the level of the latent treatment status $D_{i,g,t}^{*}$, recognizing that misclassification may occur within specific groups and periods. Furthermore, mean independence and strong exogeneity are defined with respect to the observed treatment status $\mathbf{D}_{g}$. Finally, similar to De Chaisemartin and D’Haultefoeuille (2020), we maintain a common trends assumption as our core identification assumption throughout the paper. [^3] See De Chaisemartin and D’Haultefoeuille (2020) for a discussion of the plausibility of the common trends assumption in settings with heterogeneous treatment effects.
@@
The proposition bears some resemblance to Lemma 1 of Denteh and Kedagni (2022). It shows that if misclassification is “nondifferential”, meaning it is independent of potential outcomes given the true treatment status (e.g., Lewbel, 2007a; Denteh and Kedagni, 2022), and if misclassification within each group is conditionally independent of the observed treatment statuses of other groups, then mean independence and strong exogeneity with respect to the true treatment status $\mathbf{D}_{g}^{*}$ imply the analogous assumptions with respect to the observed treatment status $\mathbf{D}_{g}$.
To further illustrate the mean independence assumption, note that (3) is equivalent to
@@
for all $d\in[0,1]$ and $d^{\prime},d^{\prime\prime}\in\{0,1\}^{T}$, provided the expectations are well-defined. This condition states that, given the actual treatment status, potential nontreatment outcomes exhibit the same trends regardless of whether the treatment status is correctly classified or misclassified. We also note that (5) is weaker than the joint independence $(\mathbf{D}_{g}^{*},\mathbf{D}_{g})\perp\mathbf{D}_{-g}$.
@@
Assumption 6 is imposed solely for simplifying the notation. For designs involving more general switching patterns, similar assumptions about misclassification can be made for different types of switchers.
@@ 3.1 Homogenous Treatment Effects
[…]
@@ 3.2 ATE of Observed Switchers
To formalize the source of bias in what follows, we focus on a specific form of misclassification that occurs in the observed untreated period(s) before switching to treatment. This type of measurement error is common in applied settings where the precise timing of treatment is not observable due to data limitations. For instance, in a study of the causal effects of capital punishment reform on jury behavior, Bindler and Hjalmarsson (2018) define the “treatment year” as the first year with zero recorded death sentences, despite acknowledging that the actual legal change may have taken place earlier. In a staggered adoption setting, such coding can lead to systematic misclassification of pre-treatment periods in treated groups. The following assumption captures this structure of misclassification around the time of observed treatment adoption:
@@
For each $(i,g,t)\in\{1,\cdots,N_{g,t}\}\times\{1,\ldots,G\}\times\{1,\ldots,T\}$ and $(j,g^{\prime},t^{\prime})\in\{1,\cdots,N_{g,t-1}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$:
@@
Assumption 7 requires that the unit-level observed treatment status is correctly classified in two situations: (i) periods in which a group is observed to receive treatment, and (ii) observed untreated periods that are followed by another observed untreated period. This assumption restricts misclassification to periods immediately preceding observed treatment, a simplification made for tractability and could be relaxed in more general settings. [^5] Note however that this formulation allows for misclassification to occur in the comparison group during the “treatment year”, which is particularly relevant when using “not-yet-treated” groups as controls. For instance, suppose a group is observed to be untreated in periods $t-1$ and $t$ (${D}_{g,t-1}=0$, ${D}_{g,t}=0$), but receives treatment in period $t+1$ (${D}_{g,t+1}=1$). Then, misclassification may arise in period $t$. In the context of Bindler and Hjalmarsson (2018), this means that units from the year before the “treatment year”, defined as the first year with zero death sentences, may be misclassified due to uncertainty about the timing of the legal reform, while units from other years are correctly classified.
Let $N_{S}:=\sum_{(g,t):t\geq 2,D_{g,t}>D_{g,t-1}}N_{g,t}$ denote the number of observations in the observed switching-to-treatment cells. In staggered adoption designs, this is the only type of observed switchers. To keep the analysis tractable, we focus on this staggered case and do not consider more general switching patterns, although such extensions are conceptually straightforward at the cost of more complex assumptions and arguments. We begin by defining a first causal parameter of interest:
@@
where the second equality uses Assumption 6. This quantity represents the ATE for units in the observed switching-to-treatment cells, where a group is observed to initiate or to receive treatment, across all switching periods among eventually treated groups.
We consider $\delta^{S}$ a natural starting point, as it coincides with the ATE of switchers in De Chaisemartin and D’Haultefoeuille (2020) when there is no misclassification. [^6] Note that, in the absence of misclassification, the ATE of switchers can in fact be viewed as a weighted average (weighted by the occurence probability of each group) of $\mathrm{ATE}(t,t)$s across time, where $\mathrm{ATE}(t,t)$ is the ATE in period $t$ of groups experiencing treatment at time $t$ (cf. Callaway and Sant’Anna, 2021). Even under misclassification, it remains a meaningful quantity, since it refers to a clearly defined subpopulation for which treatment is observed to begin in the data and serves as a useful benchmark. However, due to the patterns described in Assumption 7, some observed switching cells align with the actual switch, while others correspond to periods after treatment has already started. As a result, $\delta^{S}$ averages treatment effects from a mixture of true switchers and units who actually switched to treatment earlier than observed. While we demostrate the estimator proposed by De Chaisemartin and D’Haultefoeuille (2020) is generally biased in this case, we also show how to adjust it in the presence of misclassification. In Subsection 3.3, we also introduce the ATE of true switchers, an arguably more meaningful target of interest, which we aim to identify and estimate in Section 4.
Before introducing the estimator for $\delta^{S}$, we impose an additional assumption on the existence of comparison groups, following De Chaisemartin and D’Haultefoeuille (2020):
@@
is set to zero when there is no switching-to-treatment group ($D_{g,t}=1$, $D_{g,t-1}=0$), or no comparison group ($D_{g,t}=0,D_{g,t-1}=0$). Intuitively, $\widehat{\operatorname{DID}}$ first computes a time-specific difference-in-differences $\widehat{\operatorname{DID}}_{t}$ by comparing outcome changes between observed switching and comparison groups, and then aggregates all these two-by-two comparisons across all time periods. In staggered adoption designs, this estimator coincides with the estimator labelled as $\widehat{\mathrm{DID}}_{M}$ suggested by De Chaisemartin and D’Haultefoeuille (2020).
@@
The following proposition formalizes this argument and shows that the DID estimand, $\mathrm{E}\left[\widehat{\mathrm{DID}}\right]$ can be decomposed into three components: (1) the ATE of units in the observed switching-to-treatment cells, (2) a bias term that relates to the ATE of the misclassified, actually treated units in the observed just-before-switching cells, and (3) a second bias term that relates to the ATE of the misclassified, actually treated units in the comparison cells.
@@
$\qquad\ -\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{1,0,t}}{N_{0,0,t}}D_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right].$ (8)
The first term in (8) is the ATE of units in the observed switching-to-treatment cells. The second term captures bias arising from actually treated units in the observed just-before-switching cells, who are misclassified as untreated. It disappears if such misclassification is absent, i.e., ${D}_{i,g,t-1}^{*}=0$ for all $(i,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$ such that $D_{i,g,t-1}=0$ and $D_{i,g,t}=1$. To facilitate interpretation, assume $\frac{N_{g,t}}{N_{g,t-1}}=1$ for all $(g,t)\in\{1,\ldots,G\}\times\{2,\ldots,T\}$, a requirement that is satisfied when the data are a balanced panel. then this bias can be interpreted as the reweighted ATE of the misclassified, actually treated units in the observed just-before-switching cells,
@@
where $N_{1,1;1,0}:=\sum_{(i,g,t):t\geq 2}\mathbb{I}\left\{D_{i,g,t}^{*}=1,D_{i,g,t-1}^{*}=1,D_{i,g,t}=1,D_{i,g,t-1}=0\right\}$, and $w\left(\mathbf{D}\right):=\frac{N_{1,1;1,0}}{N_{S}}$ is the weighting factor and between zero and one with probability one. Similarly, the third is the bias due to the misclassification of the actually treated units in the comparison cells of the observed switching-to-treatment cells, which disappears if ${D}_{i,g,t}^{*}=0$ for all $(i,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$ such that $D_{i,g,t-1}=0$ and $D_{i,g,t}=0$. Loosely speaking, it relates to the ATE of the misclassified, actually treated units in comparison cells of the observed switching-to-treatment cells, however, we do not interpret it further due to the complexity of the weighting. [^7] If misclassification is not only present in periods just before observed switching-to-treatment cells, one can show that a term analogous to the third term in (8) for comparison groups appears also for period $t-1$ in the bias expression.

### v2

@@ Abstract
This paper examines the identification and estimation of treatment effects in staggered adoption designs – a common extension of the canonical Difference-in-Differences (DiD) model to multiple groups and time-periods – in the presence of (time varying) misclassification of the treatment status as well as of anticipation. We demonstrate that standard estimators are biased with respect to commonly used causal parameters of interest under such forms of misspecification. To address this issue, we provide modified estimators that recover the Average Treatment Effect of observed and true switching units, respectively. Additionally, we suggest two moment based specification tests aimed at detecting Parallel Trends violations in pre-treatment periods as well as the timing and extent of misclassification and anticipation effects. We illustrate the proposed methods with an application to the effects of an anti-cheating policy on school mean test scores in high stakes national exams in Indonesia.
@@
Keywords: Measurement Error, Causal Inference, Moment Equality Tests, Specification Tests
@@ 1 Introduction
The Difference-in-Differences (DiD) methodology is arguably one of the most popular research designs in applied economics, offering a simple and intuitive framework to estimate causal effects in observational settings. It serves as a preferred tool for policy evaluations, labor market studies, and other areas where randomized controlled trials are impractical. Staggered adoption designs naturally extend the canonical two-by-two DiD framework to settings where multiple groups adopt treatments at different times. This structure enables researchers to exploit variation in treatment timing and group-level adoption, aggregate treatment effects across groups, and accommodate heterogeneity in treatment effects. [^1] To illustrate the prevalence of staggered adoption designs, BLW2022 report that more than 54% of all DiD studies published in the leading Finance and Accounting journals between 2000 and 2019 (744 studies in total) were based on staggered adoption designs.
@@
Misclassification of the treatment status arises when the recorded treatment status does not coincide with the actual treatment status. This can occur due to data limitations, such as when the exact timing of treatment is not observed (e.g., Bindler2018), or due to reporting and classification errors when treatment status can only be inferred indirectly (see GS2017). Anticipation effects, on the other hand, occur when individuals or units are either exposed to treatment prior to the recorded treatment date or adjust their behavior in anticipation of future treatment, thereby contaminating pre-treatment periods (see, e.g., B2022; BBS2022; BPRSS2024). Both issues are particularly problematic in staggered adoption designs, where treatment assignment evolves over time and across groups, amplifying the risk of “forbidden comparisons” and leading to invalid inference, even when estimands appropriate for staggered adoption settings are targeted.
@@
As a second contribution, we define alternative estimators that allow to recover the aforementioned causal parameters in staggered adoption designs. In particular, we introduce a bias-corrected version of the staggered adoption estimator proposed by CH2020, which recovers the ATE of observed switching-to-treatment units. Based on knowledge of the timing and extent of misspecification, the estimator adjusts the comparisons conducted by the original estimator of CH2020 by using groups unaffected by misspecification. In contrast, for the ATE of actual switching-to-treatment cells, additional steps are required. Since the actual timing of treatment adoption is unobserved at the unit level, we exploit a homogeneity assumption on the misclassification (anticipation) probability, which restricts heterogeneity across units but not over time. This allows us to consistently estimate the fraction of actual switching-to-treatment units by extending the Parallel Trends (PT) assumption to the (potential) outcomes under treatment. In a subsequent step, we construct an estimator for the ATE of actual switching-to-treatment cells and derive its asymptotic properties.
As a third contribution, we propose two flexible specification tests based on moment equalities that allow to check for violations of pre-treatment PT and for the presence, extent, and timing of misclassification. In fact, testing for both forms of misspecification is required as our proposed corrections rely on PT assumptions, and distinguishing between bias arising from misclassification and bias due to PT violations in pre-treatment periods is inherently challenging. To address this, we firstly propose a moment equality based test for pre-trends. This test compares switching groups with not-yet-switched comparison groups in periods known not to be affected by misclassification or anticipation. This test may serve as a more powerful alternative to event-study type pre-trend tests as it uses all possible combinations of comparisons across the time periods considered and not just with respect to a specific reference period. The second test on the other hand is a test for misclassification or anticipation comparing the outcomes of observed switcher groups to not-yet-switched groups using periods potentially affected by the former. We implement these tests as moment equality tests and derive their asymptotic properties. Finally, we show how both tests may be combined to construct a decision rule for practicioners.
@@
To illustrate the finite sample performance of both the estimators and the tests, we carry out a small-scale Monte Carlo simulation. The findings suggest a good performance even in cases of misspecification because of misclassification or anticipation. We then apply these methods to assess the effects and the presence of anticipation in a study by BPRSS2024 on the effects of a new computer based testing system on school mean test scores in Indonesia between 2015 to 2019. The testing system, which was voluntary and intended to curb cheating behavior, was reported to have led schools to anticipate measures that limit cheating. Our estimates for the ATE of observed switching-to-treatment units confirm this and align with BPRSS2024, while the estimates for the ATE of actual switching-to-treatment units are slightly larger in magnitude. Moreover, we estimate the fraction of anticipating schools to be around 15% in 2018, and up to 30% in 2019. Finally, the misspecification tests provide evidence for anticipation effects, though they also suggest some violation of parallel pre-trends in periods just prior to that anticipation for at least one cohort.
@@
By providing a framework for inference in staggered adoption designs under misspecification, this paper contributes to the growing literature on the methodological underpinnings of DiD. In particular, the paper builds on three strands of the econometric literature. First, it contributes to the growing body of research on DiD methods tailored to staggered adoption designs, including recent advancements that address the challenges of heterogeneous treatment effects and dynamic treatment timing (e.g., CH2020; CSA2021; BJS2024). Second, the paper intersects with a broader literature on the misclassification of treatment status, which is well known to compromise the validity of inference procedures (e.g., L2007; U2018; DK2022). Unlike other approaches in the DiD literature that rely on proxy variables (e.g., BG2018) or exclusion restrictions (NN2025), our framework does not require such external information. Third, the paper relates to recent contributions on pre-testing of Parallel Trends and its implications for causal inference depending on the test outcome (e.g., Rambachan2023; CH2024).
@@ 2 Setup
We build on the setup of CH2020, adapting it to account for treatment status misclassification. Specifically, as in CH2020, we assume that the observations can be divided into $G$ groups observed over $T$ periods, with each group-period combination indexed by $(g,t)\in\{1,\ldots,G\}\times\{1,\ldots,T\}$. For each $(g,t)$ pair, we observe $N_{g,t}$ units indexed by $i\in\{1,\ldots,N_{g,t}\}$. Accordingly, for each $i$ within a given $(g,t)$ combination, we define the potential outcomes $(Y_{i,g,t}(0),Y_{i,g,t}(1))$ alongside the latent, potentially misclassified, binary treatment status $D_{i,g,t}^{*}$, where $D_{i,g,t}^{*}=1$ if unit $i$ from group $g$ receives treatment in period $t$, and $D_{i,g,t}^{*}=0$ otherwise. The observed (or recorded) treatment status is given by a separate binary indicator $D_{i,g,t}$, which may differ from $D_{i,g,t}^{*}$ with positive probability, as formalized in Assumptions 6 and 7 below. Additionally, to explicitly address potential misclassification of treatment timing under the sharp design (see Assumption 2 below), we introduce a latent indicator $S^{*}_{i,g,t}$, equal to one if unit $i$ from group $g$ actually switches to treatment at the beginning of period $t$, and zero otherwise. Unlike the observed treatment indicator $D_{i,g,t}$, which might incorrectly classify some units as switchers, $S^{*}_{i,g,t}$ precisely identifies true switching events. This distinction is especially useful in cross-sectional settings, where each unit is observed only once, allowing us to define causal parameters based explicitly on actual rather than observed switchers (see SubSection 3.3).
@@
Similar to CH2020, we replace the independent groups assumption with mean independence, since the presence of stable groups (as defined below) inherently violates independence across groups (specifically, observed treatment status cannot be independent across groups in this case). Moreover, our sharp design assumption (Assumption 2) is stated with respect to the observed treatment status $D_{i,g,t}$, thus explicitly accommodating scenarios in which researchers mistakenly assume a sharp design based solely on observed data. Importantly, we allow for violations of the sharp design at the level of the latent treatment status $D_{i,g,t}^{*}$, recognizing that misclassification may occur within specific groups and periods. Furthermore, mean independence and strong exogeneity are defined with respect to the observed treatment status $\mathbf{D}_{g}$. Finally, similar to CH2020, we maintain a common trends assumption as our core identification assumption throughout the paper. [^3] See CH2020 for a discussion of the plausibility of the common trends assumption in settings with heterogeneous treatment effects.
@@
The proposition bears some resemblance to Lemma 1 of DK2022. It shows that assumptions stated with respect to the true treatment status $\mathbf{D}^{*}$ can be translated to the observed treatment status $\mathbf{D}$ under two additional conditions. First, misclassification is trend-nondifferential in mean, in the sense that conditional on the true treatment status, the observed treatment indicator is mean-independent of the untreated outcome trend. This condition is the natural trend analogue of mean nondifferential misclassification in levels (see, e.g., L2007; DK2022). Second, misclassification is conditionally independent across groups, meaning that within a given group, the true treatment status is independent of the observed treatment paths of other groups once the group’s own observed treatment status is conditioned on.
To further clarify the mean independence assumption, note that (3) is equivalent to
@@
for all $d\in\{0,1\}$ and all $d^{\prime},d^{\prime\prime}\in\{0,1\}^{T}$, provided the expectations are well defined. This condition states that, conditional on the true treatment status, untreated potential outcomes follow the same trend regardless of whether the observed treatment status is correctly classified or misclassified. Such mean trend-nondifferential misclassification is analogous to the standard mean nondifferential misclassification assumption in the measurement error literature, and is typically plausible when misclassification arises at the data collection or recording stage. By contrast, it may fail in settings where treatment status is self-reported. Finally, we emphasize that the conditions in Proposition 1 are sufficient but not necessary for Assumptions 3 and 4 to hold. In what follows, we therefore work directly with Assumptions 3 and 4 to derive transparent bias expressions and corresponding estimators.
@@
Assumption 6 is imposed solely for simplifying the notation. For designs involving more general switching patterns, similar assumptions about misclassification can be made for different types of switchers (“reverse switchers”) at the cost of more complex notation.
@@ 3.1 Homogenous Treatment Effects
holds for all $i$, $g$, and $t$. Specifically, let $\widehat{\beta}_{fe}$ denote the TWFE estimator obtained from a regression of $Y_{i,g,t}$ on group and time fixed effects as well as the observed treatment status $D_{i,g,t}$. [^4] Under the sharp design (Assumption 2), we have $D_{i,g,t}=D_{g,t}$ for all $i\in\{1,\dots,N_{g,t}\}$ and $(g,t)$. In the absence of misclassification and under Assumptions 15, it follows from Theorem 1 of CH2020 that:
@@ 3.2 ATE of Observed Switchers
To formalize the source of bias in what follows, we focus on a specific pattern of misclassification that arises in the observed untreated period(s) prior to treatment adoption. This form of measurement error is common in applied work when the exact timing of treatment is not directly observed due to data limitations. In such cases, treatment is often coded using proxy rules that may lag the true policy change.
For example, in a study of the causal effects of capital punishment reform on jury behavior, Bindler2018 define the “treatment year” as the first year with zero recorded death sentences, while acknowledging that the underlying legal change may have occurred earlier. In a staggered adoption setting, this type of coding can generate systematic misclassification of pre-treatment periods for treated groups, even when treatment status is correctly measured once adoption is observed. The following assumption formalizes this structure of misclassification around the time of observed treatment adoption.
@@
For each $(i,g,t)\in\{1,\cdots,N_{g,t}\}\times\{1,\ldots,G\}\times\{1,\ldots,T\}$ and $(j,g^{\prime},t^{\prime})\in\{1,\cdots,N_{g^{\prime},t^{\prime}-1}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$:
@@
Assumption 7 requires that the unit-level observed treatment status is correctly classified in two situations: (i) periods in which a group is observed to receive treatment, and (ii) observed untreated periods that are followed by another observed untreated period. This assumption is a simplification made for tractability, and restricts misclassification to periods immediately preceding observed treatment. It can be relaxed in a straightforward manner to more general settings. [^5] Note however that this formulation allows for misclassification to occur in the comparison group during the “treatment year”, which is particularly relevant when using “not-yet-treated” groups as controls. For instance, suppose a group is observed to be untreated in periods t1t-1 and $t$ (${D}_{g,t-1}=0$, ${D}_{g,t}=0$), but receives treatment in period $t+1$ (${D}_{g,t+1}=1$). Then, misclassification may arise in period $t$. In the context of Bindler2018, this assumption implies that units from the year immediately preceding the “treatment year,” defined as the first year with zero death sentences, may be misclassified due to uncertainty about the timing of the legal reform, whereas units from all other years are correctly classified.
Let $N_{S}:=\sum_{(g,t):\,t\geq 2,\;D_{g,t}>D_{g,t-1}}N_{g,t}$ denote the total number of observations in observed switching-to-treatment cells. In staggered adoption designs, this is the only type of observed switching. To keep the analysis tractable, we restrict attention to this staggered setting and do not consider more general switching patterns, although such extensions are conceptually straightforward at the cost of more involved assumptions and arguments.
We begin by defining a first causal parameter of interest:
@@
where the second equality follows from Assumption 6. The parameter $\delta^{S}$ represents the average treatment effect for units in the observed switching-to-treatment cells, that is, periods in which a group is observed to initiate treatment, aggregated across all switching periods among eventually treated groups. However, under the misclassification patterns described in Assumption 7, some observed switching cells coincide with the true treatment adoption period, whereas others correspond to periods in which treatment had already been in place but is first recorded as starting in the data. Consequently, $\delta^{S}$ averages treatment effects from a mixture of true switchers and units that actually switched to treatment earlier than observed. While we show that the estimator proposed by CH2020 is generally biased in this setting, we also demonstrate how it can be adjusted to account for misclassification. Finally, in Subsection 3.3, we introduce the ATE of true switchers, an arguably more meaningful estimand, which we aim to identify and estimate in Section 4.
Before introducing the estimator for $\delta^{S}$, we impose an additional assumption on the existence of comparison groups, following CH2020:
@@
is set to zero when there is no switching-to-treatment group ($D_{g,t}=1$, $D_{g,t-1}=0$), or no comparison group ($D_{g,t}=0,D_{g,t-1}=0$). Intuitively, $\widehat{\mathrm{DID}}$ first computes a time-specific difference-in-differences estimator $\widehat{\mathrm{DID}}_{t}$ by comparing outcome changes between observed switching and comparison groups, and then aggregates these two-by-two comparisons across all periods. In staggered adoption designs, this estimator coincides with the estimator $\widehat{\mathrm{DID}}_{M}$ proposed by CH2020.
@@
The following proposition formalizes this argument by characterizing the DID estimand $\mathrm{E}\!\left[\widehat{\mathrm{DID}}\right$] and decomposing it into three components: (i) the ATE for units in the observed switching-to-treatment cells, (ii) a bias term arising from misclassified units that are already treated in the observed just-before-switching cells, and (iii) an additional bias term due to misclassified, actually treated units in the comparison cells.
@@
$\hskip 18.49988pt\ -\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{1,0,t}}{N_{0,0,t}}D_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right].$ (8)
The first term in (8) corresponds to $\delta^{S}$, the ATE for units in the observed switching-to-treatment cells. The second term captures bias arising from units that are already treated in period t1t-1 but are misclassified as untreated in the observed just-before-switching cells. It disappears if such misclassification is absent, i.e., ${D}_{i,g,t-1}^{*}=0$ for all $(i,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$ such that $D_{i,g,t-1}=0$ and $D_{i,g,t}=1$. To facilitate interpretation, assume $\frac{N_{g,t}}{N_{g,t-1}}=1$ for all $(g,t)\in\{1,\ldots,G\}\times\{2,\ldots,T\}$, a requirement that is satisfied when the data are a balanced panel. This bias can be interpreted as the reweighted ATE of the misclassified, actually treated units in the observed just-before-switching cells,
@@
where $N_{1,1;1,0}:=\sum_{(i,g,t):t\geq 2}\mathbb{I}\left\{D_{i,g,t}^{*}=1,D_{i,g,t-1}^{*}=1,D_{i,g,t}=1,D_{i,g,t-1}=0\right\}$, and $w\left(\mathbf{D}\right):=\frac{N_{1,1;1,0}}{N_{S}}$ is the weighting factor and between zero and one with probability one. The third term in (8) represents bias arising from misclassified, actually treated units in the comparison cells associated with the observed switching-to-treatment groups. This term disappears if $D_{i,g,t}^{*}=0$ for all $(i,g,t)$ such that $D_{g,t-1}=0$ and $D_{g,t}=0$. While it is also related to treatment effects for misclassified treated units, its interpretation is less transparent due to the more complex weighting structure, and we therefore do not pursue it further. [^6] If misclassification extends beyond periods immediately preceding observed switching, an additional bias term analogous to the third term in (8) arises for period t1t-1.